OpenStudy (anonymous):

Solve for x: 2 over 3 x + 1 = 27

myininaya (myininaya):

\[\frac{2}{3x+1}=27 ?\]

OpenStudy (anonymous):

I did it and got 39.

myininaya (myininaya):

if so multiply both sides by 3x+1 2=27(3x+1) distribute 2=81x+27 subtract 27 on both sides 2-27=81x -25=81x

myininaya (myininaya):

i leave the last part as an exercise to solve ax=b divide both sides by a